# Semantic Segmentation , Self-Driving Car , Computer vision , BDD100K , Light-weight model , Embedded devices

### v4

@@ 3 Proposed method
(a)Stride ESP
(b)ESP
(c)DESP
Figure 3: This figure presents the ESP blocks within the Encoder. The convolutional layers are designated as Conv-n (standard convolution), DConv-n ( dilated convolution), and DDConv-n ( depthwise dilated convolution), and are described in terms of . These blocks consolidate feature maps employing a Hierarchical Feature Fusion (HFF)[13], which is notable for its computational efficiency and its ability to eliminate grid-like artifacts resulting from dilated convolutions. The Stride ESP (a) not only accomplishes downsampling but also converts the depth of the feature map from to , whereas ESP (Fig 3(b)) and DESP (Fig 3(c)) maintain the original dimensions of the feature map.
This paper introduces TwinLiteNet+, a model specifically engineered for segmenting drivable areas and lanes. Inspired by the preceding TwinLiteNet [8], our study aims to enhance the model by refining both the encoder and decoder components. In particular, the proposed decoder is designed to effectively leverage information compared to its predecessor while still maintaining a low computational resource utilization. Moreover, the encoder integrates depth-wise dilated separable convolutions, thereby optimizing the inference time of the model which is suitable for real-time implementation. TwinLiteNet+ operates in two primary phases: the Encode-phase and the Decode-phase, incorporating a Partial Class Activation Attention (PCAA) module [33] which enhances segmentation precision and efficiency by focusing on key areas like drivable zones and lanes.
Figure 4: Comprehensive schematic of the Encoder in TwinLiteNet+.
@@ 3.1 Encoder
Figure 5: Decoder Block Design in TwinLiteNet+. This illustrates the implementation of Upper Convolution Block (UCB) and Upper Simple Block (USB) within the decoder, specifically tailored for upsampling to generate segment maps for diverse tasks.
@@ 3.2 Decoder
Table 1: Model architecture settings in different TwinLiteNet+ model’s configs: Nano, Small, Medium and Large.
Layer
Ouput Size
Output channels for TwinLiteNet+
Nano
Small
Medium
Large
Conv
4
8
16
32
Conv
8
16
32
64
Stride ESP
16
32
64
128
P DESP
16
32
64
128
Conv
32
64
128
256
Stride ESP
32
64
128
256
Q DESP
32
64
128
256
Conv
16
32
64
128
PCAA
16
32
64
128
Conv
8
16
32
64
UCB
UCB
4
8
16
32
UCB
UCB
4
8
8
8
USB
USB
2
2
2
2
P, Q
1, 1
2, 3
3, 5
5, 7
#Param.
0.03M
0.12M
0.48M
1.94M
FLOPs
0.57G
1.40G
4.63G
17.58G
@@
Algorithm 2 TwinLiteNet+ Training Stage
1:Target end-to-end network with parameters ; Training dataset ; Validation dataset
2:Well-trained network:
3:Initialize the parameters
4:for epoch: 1 epochs do
5:for each batch in do
6:
7:
8:
9:…………………..
10:
11:…………………..
12:
13:
14:
15:endfor
16:
17:
18:
19:
20:endfor
21:return
@@ 3.3 Model configurations
Our TwinLiteNet+ model is configured into four distinct configs: TwinLiteNet, TwinLiteNet, TwinLiteNet and TwinLiteNet, with TwinLiteNet being the largest config comprising 1.94M parameters, and TwinLiteNet the smallest with only 0.03M parameters. The models differ in terms of the number of kernels in convolutional layers and the hyperparameters P and Q, while retaining the overall architecture. The specifics of these configurations are detailed in Table 1. This leads to variations in the number of parameters and computational costs across the configs. The development of these four configs demonstrates our model’s versatility across different configurations and opens up options for deployment on diverse hardware platforms. This versatility is particularly significant for embedded devices, enabling the inference capabilities of the TwinLiteNet+ model, especially in the TwinLiteNet config with its 0.03M parameters and a computational requirement of 0.57 GFLOPs.
Table 2: Comparison between TwinLiteNet+ Model’s configs in FPS (in 5 different batch sizes), Parameters, FLOPs and Model size. (Parameters and FLOPs represent the number of parameters and the number of floating point operations required in each model, FPS shows the model’s latency when inference.
Config
FPS (with Batch Size)
#Param.
FLOPs(batch=1)
Model Size
1
2
4
8
16
Nano
237
470
921
1004
1163
0.03M
0.57G
0.06MB
Small
178
340
697
732
779
0.12M
1.40G
0.23MB
Medium
138
269
442
456
484
0.48M
4.63G
0.92MB
Large
109
186
218
220
237
1.94M
17.58G
3.72MB
@@ 3.4 Loss Function
In the design of our proposed segmentation model, we incorporate two distinct loss functions: Focal Loss [34] and Tversky Loss [35]. These functions are chosen to tackle specific challenges in pixel-wise classification tasks effectively.
Focal Loss [34]: Focal Loss is designed to address pixel classification errors by focusing more on misclassified samples. It modifies the standard cross-entropy loss to emphasize hard-to-predict samples, thereby improving performance in imbalanced datasets. The function is defined in Eq. 5. Where is the total number of pixels, is the number of classes, is the predicted probability of pixel for class , is the ground truth, and is a parameter that adjusts the focus on difficult samples.
Tversky Loss [35]: An adaptation of Dice Loss [38], Tversky Loss addresses class imbalances in segmentation tasks by introducing parameters and to control the impact of false positives and false negatives which is detailed in Eq. 6. where , , and are the counts of true positives, false negatives, and false positives, respectively. The parameters and allow fine-tuning the loss to emphasize precision or recall, making it a versatile tool for handling imbalanced data in segmentation tasks.
The overall loss function, which aggregates the contributions from both Focal and Tversky Losses, is formulated for each segmentation head as:
where and represent the aggregated Focal and Tversky Losses for Drivable Area and Lane segmentation, respectively.
@@ 4.1.1 Dataset
The BDD100K [39] dataset was used for training and validating TwinLiteNet+. With 100,000 frames and annotations for 10 tasks, it is a large dataset for autonomous driving. Due to its diversity in geography, environment, and weather conditions, the algorithm trained on the BDD100K dataset is robust enough to generalize to new settings. The BDD100K dataset is divided into three parts: a training set with 70,000 images, a validation set with 10,000 images, and a test set with 20,000 images. Since no labels are available for the 20,000 images in the test set, we choose to evaluate on a separate validation set of 10,000 images.
We use data that is prepared similarly to several previous studies [4, 8, 10] to ensure fairness in comparison.
@@ 4.1.2 Evaluation Metrics
For the segmentation tasks, similar to the evaluation approach in [10, 11, 29], we assess the drivable area segmentation task using the mIoU (mean Intersection of Union) metric. In the context of lane segmentation, we measure the performance using both accuracy and IoU (Intersection over Union) metrics. However, owing to pixel imbalances between the background and foreground in lane segmentation, we opt for a more meaningful balanced accuracy metric in our evaluation. Traditional accuracy metrics may produce biased results by favoring classes with a larger number of samples. In contrast, balanced accuracy provides a fairer assessment by considering the accuracy for each class. We use the accuracy metrics provided in the study [11].
All experiments used the PyTorch framework on an NVIDIA GeForce RTX A5000 GPU with 32GB of RAM and an Intel(R) Core(TM) i9-10900X processor.
Table 3: Performance benchmarking in mIoU (%) for Drivable Area Segmentation, IoU (%) and Accuracy (%) for Lane Segmentation of models with same tasks. (The number in parentheses after each model presents the following published year)
To improve performance, we apply several data augmentation techniques. To address photometric distortions, we modify the hue, saturation, and value parameters of the image. In addition, we also incorporate basic enhancement techniques to handle geometric distortions such as random translation, cropping, and horizontal flipping.
We train our model using the AdamW [44] optimizer with a learning rate of , momentum of 0.9, and weight decay of . In our TwinLiteNet+, we use EMA (Exponential Moving Average) model [45] purely as the final inference model. Additionally, we resize the original image dimensions from 1280 × 720 to 640 × 384 (TwinLiteNet+ is designed with a stride of 16 for height, so we need to resize the input image to a width of 384). For the loss function coefficients, we set = 0.7 and = 0.3 for , = 0.9 and = 0.1 for , and = 0.25 and = 2 in . All of the specified configurations are developed based on empirical evidence. Finally, we conduct training with a batch size of 16 on an RTX A5000 for 100 epochs. The model training process is described in Algorithm 2. During the model training process, we simultaneously train both tasks and employ EMA technique for weight updates following each backpropagation step. After each epoch, we evaluate the model on the validation set and focus on three key metrics: mean Intersection over Union for the Drivable Area Segmentation task (mIoUdrivable), Lane Accuracy, and Intersection over Union for the Lane Segmentation task (Accl, IoUl).
4.2 Experimental results
@@ 4.2.1 Cost Computation Performance
Groundtruth
TwinLiteNet
TwinLiteNet+Large
TwinLiteNet+Medium
TwinLiteNet+Small
TwinLiteNet+ Nano
Figure 6: Drivable Area and Lane Segmentation visualization results between TwinLiteNet+ model’s configs and TwinLiteNet.
Table 4: Perforance benchmarking in mIoU (%), PA (%) and mPA (%) between different models in Directly & Alternative Area segmentation tasks.

### v5

@@ 3 Proposed method
This paper introduces TwinLiteNet+, a model specifically engineered for segmenting drivable areas and lanes. Inspired by the preceding TwinLiteNet [12], our study aims to enhance the model by refining both the encoder and decoder components. In particular, the proposed decoder is designed to effectively leverage information compared to its predecessor while still maintaining a low computational resource utilization. Moreover, the encoder integrates depthwise dilated separable convolutions, thereby optimizing the inference time of the model which is suitable for real-time implementation. TwinLiteNet+ operates in two primary phases: the Encode-phase and the Decode-phase, incorporating a Partial Class Activation Attention (PCAA) module [26] which enhances segmentation precision and efficiency by focusing on key areas like drivable zones and lanes. While several attention mechanisms such as DANet [24], SE [25], and CBAM [27] have been proposed to enhance feature representations by modeling spatial and/or channel-wise dependencies, they typically rely on global feature statistics or self-attention across the entire feature map without explicitly incorporating semantic class information. In contrast, PCAA generates attention maps guided by partial class activation maps, which embed category-level semantics. This enables PCAA to focus selectively on regions relevant to specific classes, offering both strong discriminative capability and high computational efficiency.
@@ 3.3 Model configurations
To ensure versatility and broad applicability across diverse hardware devices, we design our TwinLiteNet+ model with four configurations: Nano, Small, Medium, and Large. These variants, denoted as TwinLiteNet, TwinLiteNet, TwinLiteNet, and TwinLiteNet, differ primarily in the number of convolutional filters and the structural hyperparameters P and Q, while preserving the overall encoder-decoder architecture. This configuration hierarchy allows for a flexible trade-off between performance and computational cost. Specifically, the Nano version contains only 0.03 million parameters and requires 0.57 GFLOPs, making it highly suitable for ultra-low-power edge devices. In contrast, the Large version reaches 1.94 million parameters and 17.58 GFLOPs, offering improved accuracy at the cost of higher resource consumption. All configurations share the same architectural stages but scale their complexity by adjusting the output channel widths at each stage. The encoder stages consist of stacked convolutional blocks with downsampling via Stride ESP and a repeated application of DESP determined by the hyperparameters P and Q, which control the depth and receptive field of the network. Table 1 provides a comprehensive overview of the architectural parameters and resource consumption across different model configurations.
This modular scaling strategy enables our model to adapt to varying deployment constraints without requiring changes in architecture design or retraining. In particular, the lightweight TwinLiteNet proves highly effective in latency-sensitive applications, while the Medium and Large configurations deliver stronger performance in scenarios with more relaxed computational budgets.
| Layer | Ouput Size | Output channels for TwinLiteNet+ | |||||
| Nano | Small | Medium | Large | ||||
| Conv | 4 | 8 | 16 | 32 | |||
| Conv | 8 | 16 | 32 | 64 | |||
| Stride ESP | 16 | 32 | 64 | 128 | |||
| P DESP | 16 | 32 | 64 | 128 | |||
| Conv | 32 | 64 | 128 | 256 | |||
| Stride ESP | 32 | 64 | 128 | 256 | |||
| Q DESP | 32 | 64 | 128 | 256 | |||
| Conv | 16 | 32 | 64 | 128 | |||
| PCAA | 16 | 32 | 64 | 128 | |||
| Conv | 8 | 16 | 32 | 64 | |||
4 | 8 | 16 | 32 | |||
4 | 8 | 8 | 8 | |||
2 | 2 | 2 | 2 | |||
| P, Q | 1, 1 | 2, 3 | 3, 5 | 5, 7 | |||
| #Param. | 0.03M | 0.12M | 0.48M | 1.94M | |||
| FLOPs | 0.57G | 1.40G | 4.63G | 17.58G | |||
@@ 3.4 Loss Function
To effectively address the specific challenges posed by pixel-wise classification in autonomous driving scenarios-particularly class imbalance and sensitivity to fine structural details-we employ a combination of Focal Loss [52] and Tversky Loss [53]. Each loss is applied independently to both segmentation heads to target complementary aspects of the task.
Focal Loss [52]: Designed to down-weight well-classified examples and focus learning on harder, misclassified pixels, Focal Loss is particularly beneficial in scenarios with significant class imbalance-such as drivable area and lane segmentation, where background pixels dominate. For example, in lane segmentation, narrow foreground structures (i.e., lanes) are often underrepresented, making them prone to being overlooked by standard losses. Focal Loss addresses this by introducing a modulating factor to the standard cross-entropy:
where, is the number of pixels, is the number of classes (foreground vs. background), is the one-hot ground truth label for pixel and class , is the predicted probability, and is a focusing parameter (commonly set to 2) that reduces the impact of easy examples.
Tversky Loss [53]: An extension of Dice Loss [56], Tversky Loss introduces tunable parameters and to balance penalties on false positives and false negatives. This is particularly important in lane segmentation, where missing thin, elongated structures is more critical than predicting false positives. The formulation is:
where , , and denote true positives, false negatives, and false positives, respectively. Choosing encourages the model to prioritize recall-an essential property for safety-critical applications like lane segmentation.
Total Loss: The total loss combines the drivable area and lane-specific losses:
Each sub-loss is a weighted combination of the Focal and Tversky losses applied to the corresponding output. Focal Loss mitigates class imbalance by encouraging the model to focus on harder pixels, while Tversky Loss enhances robustness in detecting thin structures by emphasizing recall. Together, they form a complementary loss framework that improves both the robustness and accuracy of segmentation in complex real-world driving environments.
@@ 4.1.1 Dataset
The BDD100K 111https://bdd-data.berkeley.edu dataset was used for training and validating TwinLiteNet+. With 100,000 frames and annotations for 10 tasks, it is a large dataset for autonomous driving. Due to its diversity in geography, environment, and weather conditions, the algorithm trained on the BDD100K dataset is robust enough to generalize to new settings. The BDD100K dataset is divided into three parts: a training set with 70,000 images, a validation set with 10,000 images, and a test set with 20,000 images. Since no labels are available for the 20,000 images in the test set, we choose to evaluate on a separate validation set of 10,000 images. We use data that is prepared similarly to several previous studies [6, 12] to ensure fairness in comparison.
@@ 4.1.2 Evaluation Metrics
For the segmentation tasks, similar to the evaluation approach in [12, 11, 10], we assess the drivable area segmentation task using the mIoU (mean Intersection of Union) metric. In the context of lane segmentation, we measure the performance using both accuracy and IoU (Intersection over Union) metrics. However, owing to pixel imbalances between the background and foreground in lane segmentation, we opt for a more meaningful balanced accuracy metric in our evaluation. Traditional accuracy metrics may produce biased results by favoring classes with a larger number of samples. In contrast, balanced accuracy provides a fairer assessment by considering the accuracy for each class. We use the accuracy metrics provided in the study [57]. To evaluate computational efficiency, we assess the model using two standard metrics: parameters and FLOPs. In line with established conventions in the field [58, 17], we use FLOPs as the total number of multiply-add operations needed for inference.
4.1.3 Implementation details
To improve performance, we apply several data augmentation techniques. To address photometric distortions, we modify the hue, saturation, and value parameters of the image. In addition, we also incorporate basic enhancement techniques to handle geometric distortions such as random translation, cropping, and horizontal flipping. We train our model using the AdamW [59] optimizer with a learning rate of , momentum of 0.9, and weight decay of . In our TwinLiteNet+, we use Exponential Moving Average (EMA) model [60] purely as the final inference model. Additionally, we resize the original image dimensions from 1280720 to 640384. For the loss function coefficients, we set = 0.7 and = 0.3 for , = 0.9 and = 0.1 for , and = 0.25 and = 2 in . All of the specified configurations are developed based on empirical evidence. Finally, we conduct training with a batch size of 16 on an RTX A5000 for 100 epochs. The model training and evaluation process is described in Algorithm 2. During the model training process, we simultaneously train both tasks and employ EMA technique for weight updates following each backpropagation step. After each epoch, we evaluate the model on the validation set and focus on three key metrics: mean Intersection over Union for the Drivable Area Segmentation task (mIoU), Lane Accuracy (Acc), and Intersection over Union (IoU) for the Lane Segmentation task. All experiments used the PyTorch framework on an NVIDIA GeForce RTX A5000 GPU with 32GB of RAM and an Intel(R) Core(TM) i9-10900X processor.
4.2 Main results
@@ 4.2.1 Cost Computation Performance
| Config | FPS (with Batch Size) | #Param. |
Model Size | ||||||
| 1 | 2 | 4 | 8 | 16 | ||||||
| Nano | 237 | 470 | 921 | 1004 | 1163 | 0.03M | 0.57G | 0.06MB | ||
| Small | 178 | 340 | 697 | 732 | 779 | 0.12M | 1.40G | 0.23MB | ||
| Medium | 138 | 269 | 442 | 456 | 484 | 0.48M | 4.63G | 0.92MB | ||
| Large | 109 | 186 | 218 | 220 | 237 | 1.94M | 17.58G | 3.72MB | ||
@@
| Model | Drivable Area | Lane | FLOPS | #Param. | ||
| mIoU (%) | Acc (%) | IoU (%) | ||||
| DeepLabV3+∗ [15] | 90.9 | – | 29.8 | 30.7G | 15.4M | |
| SegFormer∗ [16] | 92.3 | – | 31.7 | 12.1G | 7.2M | |
| R-CNNP [43] | 90.2 | – | 24.0 | – | – | |
| YOLOP [43] | 91.6 | – | 26.5 | 8.11G | 5.53M | |
| IALaneNet [10] | 90.5 | – | 30.4 | 89.83G | 17.05M | |
| IALaneNet [10] | 90.6 | – | 30.5 | 139.46G | 27.16M | |
| IALaneNet [10] | 91.3 | – | 31.5 | 96.52G | 18.35M | |
| IALaneNet [10] | 91.7 | – | 32.5 | 200.07G | 39.97M | |
| YOLOv8 (multi) [61] | 84.2 | 81.7 | 24.3 | – | – | |
| Sparse U-PDP [44] | 91.5 | – | 31.2 | – | – | |
| BILane [11] | 91.2 | – | 31.3 | – | 1.4M | |
| TwinLiteNet [12] | 91.3 | 77.8 | 31.1 | 3.9G | 0.44M | |
| TwinLiteNet | 87.3 | 70.2 | 23.3 | 0.57G | 0.03M | |
| TwinLiteNet | 90.6 | 75.8 | 29.3 | 1.40G | 0.12M | |
| TwinLiteNet | 92.0 | 79.1 | 32.3 | 4.63G | 0.48M | |
| TwinLiteNet | 92.9 | 81.9 | 34.2 | 17.58G | 1.94M | |